a mason can build a wall in 6 hours less than an apprentice. working together they can build a wall in 4 hours. how long would it take the apprentice working alone?
so far i have 1/x + 1/x-6 = 1/4. i think my next step is the find the lcm/lcd but that is where i am having problems.

Answer by scott8148(6628)

good so far
you don't need a lcm/lcd
just multiplying by the individual denominators will clear the fractions

multiplying by 4(x)(x-6) ___ 4(x-6) + 4x = x(x-6)

4x - 24 + 4x = x^2 - 6x

subtracting 8x-24 ___ 0 = x^2 - 14x + 24

factoring ___ 0 = (x-2)(x-12)

x = 2 and x = 12 ___ from the statement of the problem, 2 is not realistic